solve the simultaneous equations
4x+5y=0,
8x-15y=5.
x=
y=

Found 2 solutions by ewatrrr, Alan3354:
Answer by ewatrrr(24785)

 
Hi
solving the simultaneous equations
4x+5y=0
8x-15y=5 |checking our answer: 2-(-3) = 5
12x+15y=0*3
8x-15y=5
20x = 5
x = 1/4 and y = -1/5 (1 + 5y = 0, y = -1/5)
Ordered pair (1/4,-1/5) is the solution fo this system

Answer by Alan3354(69443)

4x+5y=0
8x-15y=5
---------
8x+10y=0
8x-15y=5
--------- Subtract
25y = -5
y = -1/5
------------
Sub for y and find x